OpenStudy (anonymous):

Which ratios are proportional to 32 : 40? Choose all answers that are correct. A. 4 : 5 B. 8 : 10 C. 96 : 120 D. 128 : 200 @KingOfNoOne

OpenStudy (anonymous):

i think a and c

OpenStudy (anonymous):

32/4 =8, 40/4 =10 96/3 =32 120/3 =40

OpenStudy (anonymous):

so then i would be right

OpenStudy (anonymous):

close

OpenStudy (anonymous):

three answers are correct

OpenStudy (dtan5457):

look at b again @pielover123

OpenStudy (anonymous):

oh i see now

OpenStudy (anonymous):

so a b c

OpenStudy (dtan5457):

yes sir

OpenStudy (anonymous):

thx